OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = one over fourx + 7 and that pass through the point (-8, 9). parallel: y = one over fourx - 11 perpendicular: y = -4x + 23 parallel: y = -4x + 11 perpendicular: y = one over fourx - 4 parallel: y = 4x - 7 perpendicular: y = negative one over fourx - 7 parallel: y = one over fourx + 11 perpendicular: y = -4x - 23

OpenStudy (poopsiedoodle):

y = 1/4x +7?

OpenStudy (poopsiedoodle):

I meant is that what you mean. And I think it is.

OpenStudy (anonymous):

Yes, this is what I mean.

OpenStudy (poopsiedoodle):

If it is, it's the first one, since parallel lines share slopes, and the slopes of perpendicular lines are the negative reciprocals. So, for instance, y = x + 2 is parallel to y = x + 5, and y = 2x + 2 is perpendicular to y = -x/2 + 3

OpenStudy (poopsiedoodle):

oh wait... just a second...

OpenStudy (anonymous):

ok

OpenStudy (poopsiedoodle):

y = 2x+2 would be perpendicular to y = -1/2x + 3 (from my example). But yeah, it's the first one.

OpenStudy (anonymous):

Are you positive? :-)

OpenStudy (anonymous):

parallel: y = 1/4x - 11 perpendicular: y = -4x + 23

OpenStudy (poopsiedoodle):

I'm sure :3

OpenStudy (poopsiedoodle):

Oh wait...

OpenStudy (anonymous):

Thank you for your help...

OpenStudy (poopsiedoodle):

I just noticed that they have to pass through a certain point

OpenStudy (anonymous):

ok

OpenStudy (poopsiedoodle):

Let me graph them, because it can be either A or D.

OpenStudy (anonymous):

Ok...I'll wait. ^v^

OpenStudy (poopsiedoodle):

It's the last one. Sorry about that :p

OpenStudy (anonymous):

It's fine. The answer is D.: parallel: y = 1/4x + 11 perpendicular: y = -4x - 23

OpenStudy (poopsiedoodle):

Yup :3

OpenStudy (anonymous):

Alright, I'm trusting you! Thank your for the help, would you mind helping me w/ another problem? :3

OpenStudy (poopsiedoodle):

Sure.

OpenStudy (anonymous):

Ok. Here: Alex stores fertilizer for his farm in a rectangular box. The length of the box is 6 feet, width is 3 inches, and height is 9 inches. Each bag of fertilizer contains 0.225 cubic foot of fertilizer. How many bags of fertilizer will fill the box completely? [1 foot = 12 inches]

OpenStudy (poopsiedoodle):

Oh geez, volume questions. Just a second...

OpenStudy (anonymous):

ok

OpenStudy (poopsiedoodle):

Lol, I forgot how to do one of the easiest things. Anyway, it looks like you can fit 720 bags of fertilizer in there, but that might be a tad bit off. So, let's call @OrangeMaster in here to see if I'm right.

OpenStudy (anonymous):

ok :)

OpenStudy (poopsiedoodle):

Or maybe @dmezzullo

OpenStudy (poopsiedoodle):

Seems like no one's coming >.>

OpenStudy (anonymous):

Hold on

OpenStudy (poopsiedoodle):

K. Trying to get someone in here atm.

OpenStudy (anonymous):

@e.mccormick May we have some assistance?

OpenStudy (anonymous):

@bibby may we have some help

OpenStudy (poopsiedoodle):

Yay! Bibby! :D

OpenStudy (anonymous):

@bibby is 720 correct?

OpenStudy (bibby):

Use unit cancellation.

OpenStudy (bibby):

lemme get ms paint

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nm

OpenStudy (bibby):

uhh does this help?

OpenStudy (poopsiedoodle):

bibby, the 6 is 6 feet.

OpenStudy (bibby):

kill me

OpenStudy (poopsiedoodle):

k. \kill @bibby

OpenStudy (anonymous):

So...how would I solve this problem?

OpenStudy (bibby):

fixed. still do we just divide 1944 by the number in the bottom?

OpenStudy (bibby):

I did and got 5. 5 bags are required to fill it

OpenStudy (poopsiedoodle):

Oh. Well holy crap I was way off.

OpenStudy (poopsiedoodle):

Haven't messed around with volume in a few years. Sorry about that.

OpenStudy (bibby):

Yeah. I hope you never try helping anyone again. Go back to school you pathetic nerd

OpenStudy (anonymous):

You're positive it's 5 bags?

OpenStudy (anonymous):

XD

OpenStudy (bibby):

I'll cut myself with a plastic knife if it isn't

OpenStudy (anonymous):

ooooohhhhh...that would hurt.

OpenStudy (poopsiedoodle):

@bibby pls 720 did seem like a tad bit much though. @Grama-Pimpsta, just try it :p

OpenStudy (anonymous):

ok thank you for your help

OpenStudy (bibby):

It's more important that you get the idea of unit cancellation. For ever cubic foot there are 1728 cubic inches because 12*12*12

OpenStudy (anonymous):

ok I do understand

OpenStudy (anonymous):

I will try saying the answer is 5, right? just making sure.

OpenStudy (bibby):

Don't try lol. Work out the math and make sure you get how I got my answer

OpenStudy (anonymous):

alright , don't go anywhere yet. I'm going to check it out.

OpenStudy (bibby):

You don't own me. I'pm gonna go get sour candy and water brb

OpenStudy (anonymous):

lol ok XD

OpenStudy (anonymous):

...alright, I don't see how you got 5 bags of fertilizer as your answer. Would you mind showing me how you got that step by step? :(

OpenStudy (bibby):

k. unit cancellation is something I learned in chemistry. basically you set up proportions so that you can be sure you're using the right numbers and units. for example:

OpenStudy (anonymous):

Hold on a sec, stupid thing isn't working.

OpenStudy (bibby):

yeah I'm gonna attach it in 5 seconds hold up

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Ok, so you have right now 36 inches and 576 inches.

OpenStudy (anonymous):

hello

OpenStudy (bibby):

are you talkign about your problem?

OpenStudy (bibby):

The volume is 72*3*9 because the 6 was in feet. as such I converted it to inches

OpenStudy (anonymous):

no, no. Those are the numbers you got on the attachment you sent me.

OpenStudy (anonymous):

OH 0___0 duh. Ok, 72 x 3 x 9

OpenStudy (anonymous):

That's...1,944 like you said.

OpenStudy (bibby):

Yeah those were just different examples with different numbers to demonstrate a concept. exactly. So now you know that your cube of whatever we're filling is 1944 cubic inches. Now we have to get the bag of whatever into cubic inches too

OpenStudy (anonymous):

Ok, so .225 cubic feet = ft^3 x ft^3 x ft^3 cancels out. .225 times 12?

OpenStudy (anonymous):

I get 2.7.

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@poopsiedoodle how did you do it?

OpenStudy (anonymous):

0.225 * 12 * 12 * 12 = 388.8 mate

OpenStudy (poopsiedoodle):

Probably the wrong way. 1994/2.7, except this time, I got 738. I don't know what I'm doing. Lol.

OpenStudy (anonymous):

btw I got suspended apologies.

OpenStudy (anonymous):

you were suspended for the following reason: Never tell anyone on OS that their best choice is some form of suicide #owned

OpenStudy (anonymous):

Parallel is y=1/4x + 8 Pendic. is y=4x + 41

OpenStudy (anonymous):

I know, I'm asking why @bibby got 5 bags in the second equation. :/

OpenStudy (anonymous):

That's wrong anyways.

OpenStudy (anonymous):

You divide it out. You have the original case. and you want to see how many bags itll take to completely fill it up

OpenStudy (anonymous):

oh okay just helping ight then

OpenStudy (anonymous):

How would I do it? I just need help solving it, I don't really understand. @spurdo

OpenStudy (anonymous):

Ok let's take the dimensions out of it. Let's say you have a line sized 10 units and you have a bunch of 2 unit sized pieces. How many pieces would it take to cover the line?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

Wait, all I do is divide 1,944 by 388.8 and I get 5 ok. I see how @bibby got it.

OpenStudy (anonymous):

uhhhhhh mate you gotta go back to kindergarten divide.

OpenStudy (anonymous):

5 sorry, thinking of subtraction.

OpenStudy (anonymous):

Division is technically logarithmic subtraction